# Privacy Checklist: Privacy Violation Detection Grounding on Contextual Integrity Theory

### v1

@@ Abstract
Privacy research has attracted wide attention as individuals worry that their private data can be easily leaked during interactions with smart devices, social platforms, and AI applications. Computer science researchers, on the other hand, commonly study privacy issues through privacy attacks and defenses on segmented fields. Privacy research is conducted on various sub-fields, including Computer Vision (CV), Natural Language Processing (NLP), and Computer Networks. Within each field, privacy has its own formulation. Though pioneering works on attacks and defenses reveal sensitive privacy issues, they are narrowly trapped and cannot fully cover people’s actual privacy concerns. Consequently, the research on general and human-centric privacy research remains rather unexplored. In this paper, we formulate the privacy issue as a reasoning problem rather than simple pattern matching. We ground on the Contextual Integrity (CI) theory which posits that people’s perceptions of privacy are highly correlated with the corresponding social context. Based on such an assumption, we develop the first comprehensive checklist that covers social identities, private attributes, and existing privacy regulations. Unlike prior works on CI that either cover limited expert annotated norms or model incomplete social context, our proposed privacy checklist uses the whole Health Insurance Portability and Accountability Act of 1996 (HIPAA) as an example, to show that we can resort to large language models (LLMs) to completely cover the HIPAA’s regulations. Additionally, our checklist also gathers expert annotations across multiple ontologies to determine private information including but not limited to personally identifiable information (PII). We use our preliminary results on the HIPAA to shed light on future context-centric privacy research to cover more privacy regulations, social norms and standards.
Introduction
@@
As deep learning models and their applications achieve wide accessibility and applicability, they revolutionize our daily lives. On the one hand, these data-driven models achieve consistent utility improvements with increasing training corpus. On the other hand, private data may be collected unintentionally, which leads to individuals’ increasing privacy concerns. To achieve trustworthy AI, great efforts have been made to study privacy issues with regard to emerging applications (Li et al. 2023; Debenedetti et al. 2023).
@@
Existing privacy studies conducted by computer science researchers are fragmented. Each field has its own privacy formulation based on specific tasks. For example, in computer networks, network communications’ confidentiality is widely discussed (Nath Nayak and Ghosh Samaddar 2010; Fonseca, Vieira, and Madeira 2009) and cryptographic methods are proposed to protect cybersecurity (Buehrer, Weide, and Sivilotti 2005; Needham and Schroeder 1978). When it comes to deep learning, in both Computer Vision and Natural Language Processing fields, there is a significant body of research on privacy attacks (Shokri et al. 2016; Carlini et al. 2021; Zhu, Liu, and Han 2019; Li, Xu, and Song 2023) and defenses (Abadi et al. 2016).
@@
Figure 1: The exemplary case to transform privacy issues into reasoning problems.
@@
In addition to the fragmented formulations, the current scope of privacy research is narrow. Most existing privacy studies focus on protecting private attributes specified by privacy regulations. Though these sensitive attributes are indeed crucial components of privacy regulations, they only constitute a small fraction of regulations’ coverage. However, privacy is not solely about protecting attributes. Instead, It involves understanding the context to ensure compliance with existing regulations and norms. For example, according to the Health Insurance Portability and Accountability Act of 1996 (HIPAA), protected health information (PHI) is considered sensitive and should be protected. However, as shown in Figure 1, though Bob’s surgery operative report is classified as protected health information, if we identify Alice as a covered entity and Bob as the individual (patient), we may apply HIPAA 164.502(a)(1)(i) to permit surgeon Alice to share Bob’s PHI with him. This example demonstrates that privacy issues can be resolved through reasoning. When we consider privacy, we actually evaluate whether our information transmission contexts violate existing standards and expectations.
@@
To bridge the gap, Contextual Integrity (CI) theory (Nissenbaum 2010) is proposed to examine the appropriateness of privacy information flows based on specific contexts. A few previous works on CI resort to formal languages such as first-order logic (Barth et al. 2006; Benthall, Gürses, and Nissenbaum 2017; Shvartzshnaider et al. 2016) to fit contexts into pre-defined clauses to determine privacy violations. However, it requires heavy effort for experts to annotate complicated legal documents into logical language. Consequently, none of these approaches can scale up to cover even a whole document.
@@
In this paper, to improve the scalability, we present Privacy Checklist. We formulate the privacy evaluation as an in-context reasoning problem with existing large language models (LLMs). Instead of annotating separate clauses, our Privacy Checklist uses large language models to extract CI characteristics and use the tree structure to capture their hierarchical structure. In addition, our Privacy Checklist also includes role-based and attribute-based graphs, as well as a definition dictionary to facilitate in-context reasoning. To determine privacy violations for the given context, we first fetch relevant regulations and use large language models to conduct in-context reasoning with retrieved regulation. We annotate all information transmission norms inside the HIPAA and show that our Privacy Checklist can improve 6% - 18% accuracy on average for existing LLMs to improve their privacy judgment ability in real court cases. Our contribution can be summarized as follows:
@@
1) We extend prior works on CI to natural language and formulate the privacy research as an in-context reasoning problem with the help of large language models.
@@
2) We propose Privacy Checklist, a first scalable knowledge base that can cover all norms of the HIPAA.
@@
3) We consider various retrieval-augmented generation (RAG) pipelines for LLMs. To retrieve relevant legal documents, we implement term frequency, semantic similarity, and agent-based methodologies.
@@
4) We conduct comprehensive experiments to demonstrate that our Privacy Checklist is effective in improving LLMs’ privacy judgment ability for real court cases.
Related Works
Privacy leakage on deep learning models.
Recent studies on privacy issues can be categorized as training data leakage and inference data leakage. For training data leakage, private data may be unintentionally collected for training and extracted via training data extraction attacks (Carlini et al. 2021; Li et al. 2023; Deng et al. 2023). In terms of inference data leakage, the users’ private data used during the inference stage may also be recovered. For example, sensitive system-level prompts may be extracted via prompt injection attacks (Wan et al. 2023; Perez and Ribeiro 2022; Liu et al. 2023; Shu et al. 2023) and side-channel attacks (Debenedetti et al. 2023). Moreover, information leakage (Gu et al. 2023; Li, Song, and Fan 2022; Song and Raghunathan 2020) may also occur on models’ hidden representation to recover sensitive information.
@@
Contextual Integrity (Nissenbaum 2010), developed by philosopher Helen Nissenbaum, posits that privacy is related to information flows and information flows must adhere to the established norms of the context to protect privacy. Each norm can typically be viewed as an atomic unit of information transmission involving three key actors: the sender, the receiver, and the data subject. Depending on the social context of these actors, such as their roles and the disclosed attributes of the data subject, we can determine whether the norm is positive or negative, which means the norm is either permitted or prohibited by existing privacy regulations. Previous works on CI (Barth et al. 2006; Benthall, Gürses, and Nissenbaum 2017; Shvartzshnaider et al. 2016) aimed to transform the context into explicit formal languages, such as first-order logic, to determine privacy violations with rule-based frameworks. However, as shown in Figure 2 (a), it costs heavy annotation efforts for experts to annotate such a knowledge base of norms. Experts must meticulously define a set of predicates and annotate the entire body of regulations, subpart by subpart. Consequently, these works on formal languages fail to scale up their norms due to the inherent complexity of legal documents. In addition, existing legal LLMs, including LawGPT (Zhou et al. 2024), Lawyer LLaMA (Huang et al. 2023) and ChatLaw (Cui et al. 2023) are proposed on the general legal domain and underperform on the legal domain due to data scarcity.
Privacy Checklist
Figure 2: The overview of privacy reasoning within the given contexts. Subfigure (a) illustrates previous approaches that use formal languages to determine privacy violations based on rules of inference and axioms. Instead, in subfigure (b), we propose an in-context reasoning pipeline with our proposed Privacy Checklist and large language models.
In this section, we show how to construct our Privacy Checklist on existing privacy regulations with less annotation cost. For each regulation, our Privacy Checklist encompasses a document tree with CI characteristics, a role graph, an attribute graph and a definition dictionary.
Legal Document Processing
@@
Due to the complexity of legal documents, each leaf node’s specification can either be a regulatory norm or clarification about new definitions. Such clarifications are irrelevant to information transmission but play crucial roles in regulatory norms. Therefore, for each specification, we ask GPT-4 to conduct a three-way classification to determine if the specification is a positive norm, negative norm, or general definition. A positive norm denotes the information transmission action (norm) that is permitted by law, while a negative norm prohibits such information transmission. Otherwise, the specification is regarded as the general definition, which enhances the norms’ clarity.
@@
Previous works, as shown in Figure 2 (a), commonly treat each norm as an isolated individual and discard the cross-references among norms. However, legal regulations are rather complicated and referring to these references is essential to assess privacy violations accurately. For instance, HIPAA 164.502(a)(1)(iv) states that “A covered entity is permitted to use or disclose protected health information except for uses and disclosures prohibited under 164.502(a)(5)(i), pursuant to and in compliance with a valid authorization under 164.508.” To permit information transmission with 164.502(a)(1)(iv), we must first ensure that no restrictions apply under 164.502(a)(5)(i) and then confirm that the authorization (one type of consent form) is valid under 164.508. Therefore, we also extend our checklist to include the annotated references for all norms in $\mathcal{T}$. For each norm, we prompt its content to GPT-4 and ask about the relationship between references inside the node attribute and the specified norm. Each reference is then categorized as either a support or an exception of the given norm.
Auxiliary Knowledge Collection
@@
Role instances extracted from daily context frequently misalign with the abstract roles defined in regulations. For example, in HIPAA, the term covered entity may refer to 1) a health plan, 2) a health care clearinghouse or 3) a health care provider. To permit Alice’s action in Figure 1, we need to link Alice’s role, surgeon, to the covered entity. Based on this observation, we further curate a role graph $\mathcal{G}^{r}=\{\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{R}^{r}\}$ where vertices $v\in\mathcal{V}^{r}$ represent social roles and only two relations “subsume” and “is subsumed by” are in $\mathcal{R}^{r}$. To construct our role graph $\mathcal{G}^{r}$, we align the roles in the legal documents with the WordNet (Miller 1994). We first use GPT-4 and parsers to recognize social roles in the legal document, then we align these roles in the WordNet and recursively append their hypernyms until the role root “person.n.01” is reached. For new definitions of roles such as the covered entity, if these definitions include roles in $\mathcal{V}^{r}$, we further append these new definitions as roles.
@@
Similarly, to identify whether the transmitted attributes are sensitive, we need an attribute graph $\mathcal{G}^{a}=\{\mathcal{V}^{a},\mathcal{E}^{a},\mathcal{R}^{a}\}$ where vertices $v\in\mathcal{V}^{a}$ represent private attributes and only two relations “subsume” and “is subsumed by” are in $\mathcal{R}^{r}$. With $\mathcal{G}^{a}$, we can map the surgery operative report to the protected health information (PHI) specified by HIPAA. To collect private attributes protected by existing regulations, we incorporate expert annotated ontologies, including Data Privacy Vocabulary (Pandit et al. 2019) and OPPO (Gupta and Hahmann 2023) into $\mathcal{G}^{a}$. Our $\mathcal{G}^{a}$ gathers the ontologies’ inclusion relationships for both class-class and class-individual. Then, for new terms that include attributes in $\mathcal{V}^{a}$, we further append them into $\mathcal{G}^{a}$.
@@
Data Summary and Manual Validation
@@
To ensure the quality and correctness of our Privacy Checklist, two of the authors manually validate positive and negative norms with the original regulations and update the inconsistently parsed CI characteristics.
Enhance In-context Reasoning with Checklist
@@
Context Annotation
In addition to directly feeding the whole context for retrieval, we adopt the following annotation methods:
@@
We follow the same CI characteristics extraction procedures used for building our Privacy Checklist to extract the sender role, receiver role, subject role, transmitted attribute, purpose, and consent form from the given context.
@@
BM25 retrieves relevant documents based on the term frequency-inverse document frequency (TF-IDF) score (Salton and Buckley 1988) and word frequency. In this context, we use the LLM explanation as the query $Q$, aiming to select the most relevant regulations by measuring the similarity between the LLM explanation query $Q$ and the regulation texts $E$ in the document tree $\mathcal{T}$.
For a word $w$ in the query $Q$, the relationship between the word and the regulation definition is calculated as follows:
@@
where ${\rm IDF}(w)$ represents the inverse document frequency of the word $w$, $f(w,E)$ denotes the frequency of the word $w$ appearing in $E$, $|E|$ is the length of $E$, and ${\rm avgdl}$ is the average length across the subrules in $\mathcal{T}$. $k_{1}$ and $b$ are hyperparameters, and we set $k_{1}=1.5$ and $b=0.75$.
The similarity between a query $Q$ and the definition $E$ is then calculated by summing the scores of all the words in $Q$:
@@
The regulations with the highest scores are then selected.
@@
In addition to statistical similarity, we also implement retrieval methods based on semantic similarity. Specifically, we use pretrained Sentence Transformers (Reimers and Gurevych 2019) to calculate both contexts and regulations’ sentence embeddings. Then, the cosine similarities are calculated for embeddings of the contexts and regulations. Notably, besides calculating semantic similarity using chunks of text, we can also calculate semantic similarity on roles and attributes. These roles and attributes can then be used as keywords to retrieve applicable rules. For simplicity, we use “all-mpnet-base-v2” (Song et al. 2020) as our embedding model throughout the paper.
@@
Besides directly retrieving relevant rules from our Privacy Checklist, we consider LLMs themselves as knowledge bases (Petroni et al. 2019) that have preliminary legal knowledge. Based on this heuristic, we implement agent-based retrieval by treating powerful LLMs as lawyers and use prompt engineering to generate applicable rules with corresponding IDs. To mitigate misinformation and hallucination, the agent-based retrieval further leverages our Privacy Checklist to verify if the generated IDs are valid.
In-context Reasoning
With the retrieved rules, we may leverage existing LLMs to perform in-context reasoning. For each candidate rule, we apply a Chain-of-Thought (CoT) prompting template (Wei et al. 2022; Wang et al. 2023) to assess its applicability in judging privacy violations within a given context. Our CoT prompt template includes four components. First, the LLM is instructed to identify the information transmission flow and analyze the three stakeholders, as well as the transmitted attributes and purpose from the given context. Second, the LLM reviews the analyzed information and determines if it matches the candidate rule’s legal terms. Third, the LLM assesses whether the candidate rule is relevant to the context. Lastly, the LLM decides whether the context is in compliance, in violation, or unrelated to the candidate rule.
Experimental Setups

### v2

@@ Abstract
Privacy research has attracted wide attention as individuals worry that their private data can be easily leaked during interactions with smart devices, social platforms, and AI applications. Existing works mostly consider privacy attacks and defenses on various sub-fields. Within each field, various privacy attacks and defenses are studied to address patterns of personally identifiable information (PII). In this paper, we argue that privacy is not solely about PII patterns. We ground on the Contextual Integrity (CI) theory which posits that people’s perceptions of privacy are highly correlated with the corresponding social context. Based on such an assumption, we formulate privacy as a reasoning problem rather than naive PII matching. We develop the first comprehensive checklist that covers social identities, private attributes, and existing privacy regulations. Unlike prior works on CI that either cover limited expert annotated norms or model incomplete social context, our proposed privacy checklist uses the whole Health Insurance Portability and Accountability Act of 1996 (HIPAA) as an example, to show that we can resort to large language models (LLMs) to completely cover the HIPAA’s regulations. Additionally, our checklist also gathers expert annotations across multiple ontologies to determine private information including but not limited to PII. We use our preliminary results on the HIPAA to shed light on future context-centric privacy research to cover more privacy regulations, social norms and standards.
Privacy Checklist: Privacy Violation Detection Grounding on Contextual Integrity Theory
Haoran Li1thanks: Haoran Li and Wei Fan contributed equally., Wei Fan [^1] footnotemark: 1, Yulin Chen2, Jiayang Cheng1 Tianshu Chu3, Xuebing Zhou3, Peizhao Hu3, Yangqiu Song1 1HKUST, 2National University of Singapore, 3Huawei Technologies {hlibt, wfanag, jchengaj}@connect.ust.hk, chenyulin28@u.nus.edu {chutianshu3, Xuebing.Zhou, hu.peizhao}@huawei.com, yqsong@cse.ust.hk
1 Introduction
@@
As machine learning models and their applications achieve wide accessibility and applicability, they revolutionize our daily lives. On the one hand, these data-driven models achieve consistent utility improvements with increasing training corpus. On the other hand, private data may be collected unintentionally, which leads to individuals’ increasing privacy concerns. To achieve trustworthy AI, great efforts have been made to study privacy issues with regard to emerging applications Li et al. (2023a); Debenedetti et al. (2023).
@@
Existing privacy studies conducted by researchers are relatively fragmented. Each field has its own privacy formulation based on specific tasks. For example, in computer networks, network communications’ confidentiality is widely discussed Nath Nayak and Ghosh Samaddar (2010); Fonseca et al. (2009) and cryptographic methods are proposed to protect cybersecurity Buehrer et al. (2005); Needham and Schroeder (1978). When it comes to deep learning, in both Computer Vision and Natural Language Processing fields, there is a significant body of research on privacy attacks Shokri et al. (2016); Carlini et al. (2021); Li et al. (2023b) and defenses Abadi et al. (2016).
@@
Figure 1: The exemplary case to transform privacy issues into reasoning problems. Our proposed checklist collects roles, attributes, transmission principles and annotated legal norms to facilitate the reasoning process.
@@
In addition to the fragmented formulations, the current scope of privacy research is narrow. Most existing privacy studies focus on protecting private attributes specified by privacy regulations. Though these sensitive attributes are indeed crucial components of privacy regulations, they only constitute a small fraction of regulations’ coverage. However, privacy is not solely about protecting attributes. Instead, It involves understanding the context to ensure compliance with existing regulations and norms. For example, according to the Health Insurance Portability and Accountability Act of 1996 (HIPAA), protected health information (PHI) is considered sensitive and should be protected. However, as shown in Figure 1, though Bob’s surgery operative report is classified as protected health information, if we identify Alice as a covered entity and Bob as the individual (patient), we may apply HIPAA 164.502(a)(1)(i) to permit surgeon Alice to share Bob’s PHI with him. This example demonstrates that privacy issues can be resolved through reasoning within the context. When we consider privacy, we actually evaluate whether our information transmission contexts violate existing standards and expectations.
@@
To bridge the gap, Contextual Integrity (CI) theory Nissenbaum (2010) is proposed to examine the appropriateness of privacy information flows based on specific contexts. A few prior works on CI resort to formal languages such as first-order logic Barth et al. (2006); Benthall et al. (2017); Shvartzshnaider et al. (2016) to fit contexts into pre-defined clauses to determine privacy violations. However, it requires heavy effort for experts to annotate complicated legal documents into logical language. Consequently, none of these approaches can scale up to cover even a whole document.
@@
In this paper, to leverage the reasoning capability of privacy information flows while allowing substantial flexibility to handle natural language ambiguity and variability, we present Privacy Checklist. We formulate the privacy evaluation as an in-context reasoning problem with existing large language models (LLMs). Instead of annotating separate clauses, our Privacy Checklist uses large language models to extract CI characteristics and use the tree structure to capture their hierarchical nature. In addition, our Privacy Checklist also includes role-based and attribute-based graphs, as well as a definition dictionary to facilitate in-context reasoning. To determine privacy violations for the given context, we first fetch relevant regulations and use large language models to conduct in-context reasoning with retrieved regulation. We annotate all information transmission norms inside the HIPAA and show that our Privacy Checklist can improve 6% - 18% accuracy on average for existing LLMs to improve their privacy judgment ability in real court cases. Our contribution can be summarized as follows [^1] Code is publicly available at https://github.com/HKUST-KnowComp/privacy_checklist.:
@@
1) We extend prior works on CI to natural language and formulate the privacy issue as an in-context reasoning problem with the help of LLMs.
@@
2) We present Privacy Checklist, a first scalable knowledge base that can cover all norms of information transmission inside the HIPAA.
@@
3) We consider various retrieval-augmented generation (RAG) pipelines. To retrieve relevant legal documents, we implement term frequency, semantic similarity, and agent-based methodologies.
@@
4) We conduct comprehensive experiments to show that our checklist is effective in improving LLMs’ privacy judgment for court cases.
2 Preliminaries
@@
Contextual Integrity Nissenbaum (2010) posits that privacy is related to information flows and information flows must adhere to the established norms of the context to protect privacy. Each norm can typically be viewed as an atomic unit of information transmission involving three key actors: the sender, the receiver, and the data subject. Depending on the social context of these actors, such as their roles and the disclosed attributes of the data subject, we can determine whether the norm is positive or negative, which means the norm is either permitted or prohibited by existing privacy regulations. Previous works on CI Barth et al. (2006); Benthall et al. (2017); Shvartzshnaider et al. (2016) aimed to transform the context into explicit formal languages, such as first-order logic, to determine privacy violations with rule-based frameworks. However, as shown in Figure 2 (a), it costs heavily for experts to annotate such a knowledge base of norms. Experts must meticulously define a set of predicates and annotate the entire body of regulations, subpart by subpart. Consequently, these works on formal languages fail to scale up their norms due to the inherent complexity of legal documents. On the other hand, existing legal LLMs, including LawGPT Zhou et al. (2024), Lawyer LLaMA Huang et al. (2023) and ChatLaw Cui et al. (2023) are proposed on the general legal domain and underperform on the legal domain due to data scarcity. GoldCoin Fan et al. (2024) proposed instruction tuning grounded on CI to enhance LLMs’ judgment ability. Instead, our work reveals the flaws of synthetic data and shows that the proper RAG pipelines without further tuning can yield comparable or even better performance.
In addition, for discussions about current privacy attacks and defenses, please refer to Appendix C.
3 Privacy Checklist Construction
Figure 2: The overview of privacy reasoning within the given contexts. Subfigure (a) illustrates previous approaches that use formal languages to determine privacy violations based on rules of inference and axioms. Instead, in subfigure (b), we propose an in-context reasoning pipeline with our proposed Privacy Checklist and LLMs.
In this section, we show how to construct our Privacy Checklist based on existing privacy regulations with less annotation cost. For the given regulation, our Privacy Checklist encompasses a document tree with CI characteristics, a role graph, an attribute graph, and a definition dictionary.
3.1 Legal Document Processing
@@ 3.2 CI Characteristics Annotation
Due to the complexity of legal documents, each leaf node’s specification can either be a regulatory norm or a clarification of the norm. Such clarifications are irrelevant to information transmission but play crucial roles in regulatory norms. Therefore, for each specification, we ask GPT-4 to conduct a three-way classification to determine if the specification is a positive norm, negative norm, or general definition. A positive norm denotes the information transmission action (norm) that is permitted by law, while a negative norm prohibits such information transmission. Otherwise, the specification is regarded as the general definition, which enhances the norms’ clarity.
@@
Previous works, as shown in Figure 2 (a), commonly treat each norm as an isolated individual and discard the cross-references among norms. However, legal regulations are rather complicated, and referring to these references is essential to assessing privacy violations accurately. For instance, HIPAA 164.502(a)(1)(iv) states that “A covered entity is permitted to use or disclose protected health information except for uses and disclosures prohibited under 164.502(a)(5)(i), pursuant to and in compliance with a valid authorization under 164.508.” To permit data transmission with 164.502(a)(1)(iv), we must first ensure that no restrictions apply under 164.502(a)(5)(i) and then confirm that the authorization (one type of consent form) is valid under 164.508. Therefore, we also extend our checklist to include the annotated references for all norms in $\mathcal{T}$. For each norm, we prompt its content to GPT-4 and ask about the relationship between references inside the node attribute and the specified norm. Each reference is then categorized as either a support or an exception of the given norm.
3.3 Auxiliary Knowledge Collection
@@
Role instances extracted from daily context frequently misalign with the abstract roles defined in regulations. For example, in HIPAA, the term covered entity may refer to 1) a health plan, 2) a health care clearinghouse or 3) a health care provider. To permit Alice’s action in Figure 1, we need to link Alice’s role, surgeon, to the covered entity. Based on this observation, we further curate a role graph $\mathcal{G}^{r}=\{\mathcal{V}^{r},\mathcal{E}^{r},\mathcal{R}^{r}\}$ where vertices $v\in\mathcal{V}^{r}$ represent social roles and only two relations “subsume” and “is subsumed by” are in $\mathcal{R}^{r}$. To construct our role graph $\mathcal{G}^{r}$, we align the roles in the legal documents with the WordNet Miller (1994). We first use GPT-4 and parsers to recognize social roles in the legal document, then we align these roles in the WordNet and recursively append their hypernyms until the role root “person.n.01” is reached. For new definitions of roles such as the covered entity, if these definitions include roles in $\mathcal{V}^{r}$, we further append these new definitions as roles.
@@
Similarly, to identify whether the transmitted attributes are sensitive, we need an attribute graph $\mathcal{G}^{a}=\{\mathcal{V}^{a},\mathcal{E}^{a},\mathcal{R}^{a}\}$ where vertices $v\in\mathcal{V}^{a}$ represent private attributes and only two relations “subsume” and “is subsumed by” are in $\mathcal{R}^{r}$. With $\mathcal{G}^{a}$, we can map the surgery operative report to the protected health information (PHI) specified by HIPAA. To collect private attributes protected by existing regulations, we incorporate expert annotated ontologies, including Data Privacy Vocabulary Pandit et al. (2019) and OPPO Gupta and Hahmann (2023) into $\mathcal{G}^{a}$. Our $\mathcal{G}^{a}$ gathers the ontologies’ inclusion relationships for both class-class and class-individual. Then, for new terms that include attributes in $\mathcal{V}^{a}$, we further append them into $\mathcal{G}^{a}$.
@@
3.4 Data Summary and Manual Validation
@@
To ensure the quality and correctness of our Privacy Checklist, two authors and one law student manually validate positive and negative norms with the original regulations and rectify the inconsistently parsed CI characteristics.
4 Enhance Reasoning with Checklist
@@
4.1 Context Annotation Methods
In addition to directly feeding the whole context for retrieval, we adopt the following annotation methods for a given context:
@@ 4.1 Context Annotation Methods
We follow the same CI characteristics extraction procedures used for building our checklist to extract the sender role, receiver role, subject role, transmitted attribute, purpose, and consent form from the given context.
@@ 4.2 Rule Retrieval Methods
BM25 retrieves documents based on the term frequency-inverse document frequency (TF-IDF) score Salton and Buckley (1988) and word frequency. We use the LLM explanation as the query $Q$, aiming to select the most relevant regulations by measuring the similarity between the LLM explanation query $Q$ and the regulation texts $E$ in the document tree $\mathcal{T}$. For a word $w$ in the query $Q$, the relationship between the word and the regulation definition is calculated as follows:
@@
where ${\rm IDF}(w)$ represents the inverse document frequency of the word $w$, $f(w,E)$ denotes the frequency of the word $w$ appearing in $E$, $|E|$ is the length of $E$, and ${\rm avgdl}$ is the average length across the subrules in $\mathcal{T}$. $k_{1}$ and $b$ are hyperparameters, and we set $k_{1}=1.5$ and $b=0.75$. The similarity between a query $Q$ and the definition $E$ is calculated by summing the scores of words in $Q$:
@@
The regulations with the highest scores are then selected as the retrieved results.
@@
In addition to statistical similarity, we also implement retrieval methods based on semantic similarity. Specifically, we use pretrained Sentence Transformers Reimers and Gurevych (2019) to calculate both contexts and regulations’ embeddings. Then, the cosine similarities are calculated for embeddings of the contexts and regulations. Notably, besides calculating semantic similarity using chunks of text, we can also calculate semantic similarity on roles and attributes. These roles and attributes can then be used as keywords to retrieve applicable rules. For simplicity, we use “all-mpnet-base-v2” Song et al. (2020) as our embedding model throughout the paper.
@@
Besides directly retrieving relevant rules from our Privacy Checklist, we consider LLMs themselves as knowledge bases Petroni et al. (2019) that have preliminary legal knowledge. Based on this heuristic, we implement agent-based retrieval by treating powerful LLMs as lawyers and use prompt engineering to generate applicable rules with corresponding IDs. To mitigate misinformation and hallucination, the agent-based retrieval further leverages our Privacy Checklist to verify if the generated IDs are valid.
4.3 In-context Reasoning
With the retrieved rules, we may leverage existing LLMs to perform in-context reasoning. For each candidate rule, we apply a Chain-of-Thought (CoT) prompting template Wei et al. (2022); Wang et al. (2023) to assess its applicability in judging privacy violations within a given context. Our CoT prompt template includes four components. First, the LLM is instructed to identify the information transmission flow and analyze the three stakeholders, as well as the transmitted attributes and purpose from the given context. Second, the LLM reviews the analyzed information and determines if it matches the candidate rule’s legal terms. Third, the LLM assesses whether the candidate rule is relevant to the context. Lastly, the LLM decides whether the context is in compliance, in violation, or unrelated to the candidate rule.
5 Experimental Setups